McDormand, Inc., reported a $3,400 unfavorable price variance for variable overhead and a $34,000 unfavorable price variancePrice variance Production volume variance Fixed overhead cost variance < Required A Required B
Answer #1
Variable overhead price variance = 3400 U
actual input at standard rate = (34100*30) = 1023000
standard rate = 1083000/36100 =30
variable overhead efficiency variance = (1083000-1023000)
variable overhead efficiency variance = 60000 F
Variable overhead cost variance = (60000-3400 ) = 56600
b) Fixed overhead price variance = 34000 U
actual input at standard rate = 34100*30 = 1023000
fixed overhead actual 1810400-(1023000+3400) = 784000
budgeted fixed overhead = (784000-34000) = 750000
fixed overhead applied (750000/37500)*36100 = 722000
Production volume variance = (750000-722000) = 28000 U
Fixed overhead cost variance = (28000+34000) = 62000 U